# Witnessing Disorder in Quantum Magnets

### v2

@@ Abstract
There are no clean samples in nature. Therefore, when we come to discuss the entanglement properties of quantum materials, the effects of disorder must be taken into account. This question is of particular interest for high-entangled states, such as quantum spin liquids, which lie outside the Landau paradigm for classifying phases of matter. In this work, we explore what experimentally-accessible measures, in the form of concurrence, residual tangle and quantum Fisher information, can teach us about the entanglement in the presence of disorder. As a representative example, we consider the Tomonaga-Luttinger liquids (TLL) and disorder-driven random singlet (RS) states found in antiferromagnetic quantum spin chains. Using quantum Fisher information and residual tangle, we demonstrate that both TLL and RS states exhibit multi-partite entanglement. In the case of the RS state, we attribute this to entanglement localized below a crossover length scale. We further show that the order of disorder average matters in calculating measures like concurrence, and that this can lead to false inferences when interpreting experiment. None the less, correctly interpreted, these witnesses provide useful information about the effects of disorder. We explore how information about the central charge of the TLL can be extracted from the low-temperature behavior of concurrence, and conjecture that this analysis can be extended to the effective central charge of the RS state. Finally, we establish how RS and TLL states can be distinguished through the growth of multi-partite entanglement, as witnessed by the equal-time structure factor. These results establish that, used carefully, experiments based on entanglement witnesses can provide important information about quantum spin systems in the presence of disorder.
@@ I Introduction
How does disorder affect entanglement within quantum materials? Can we quantify such effects experimentally? In nature, all materials possess a certain degree of disorder. Thus, when trying to understand any phase of matter, the role of disorder is an important factor to take into consideration. Disorder or randomness is especially problematic when trying to identify quantum spin liquid (QSL) states within candidate materials. This question arises in the context of many different materials, including $\kappa$-ET2Cu2(CN)3 [1, 2, 3, 4, 5, 6, 7, 8] , ZnCu3(OD)6Cl2 (Herbertsmithite) [9, 10, 11, 12, 13, 14, 7, 15, 16, 17, 18], YbMgGaO4 [19, 20, 21, 22], H3LiIr2O6 [23, 24, 25], etc. Here the challenge lies in being able to distinguish quantum fluctuations from disorder-induced ones. This puts into question whether it is the QSL or some disorder induced state realized within these materials?
Entanglement is a valuable resource to identify states such as QSLs [26]. Intuitively, one expects disorder, which localizes correlations, to have an effect on entanglement. The key issue, however, is whether this information can be accessed experimentally. Unlike cold atom systems where non local measures like (Rényi) entanglement entropy are accessible [27, 28], the same does not extend to condensed matter systems. Consequently, only limited information about the underlying state can typically be accessed experimentally. It is thus important to determine if useful information about the entanglement structure can be obtained within experimental constraints. Can we do so by making use of experimentally-accessible entanglement measure [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59], for instance?
In this study, we explore how measures based on entanglement witnesses can be used to distinguish the different phases which arise in quantum spin systems with and without disorder. As a representative example, we consider the $S=1/2$ Heisenberg (XXX) spin chain with and without bond disorder. The competing phases in this case are the Tomonaga Luttinger liquid (TLL) [60] found in the absence of disorder, and the random singlet (RS) state realised for arbitrarily weak bond disorder [61, 62, 63, 64, 65]. We consider three different measures, namely concurrence [66, 67], tangle [68] and quantum Fisher information (QFI) [40, 48], all of which have previously been discussed in the context of experiment [53, 55, 57, 59], and use exact diagonalization to explore how each of these behaves in the TLL and RS states.
From this analysis, we conclude:
  • Both the TLL and RS state are multi-partite entangled. In the TLL, multi-partite entanglement is present at all length scales. However, in the RS state, multi-partite entanglement is restricted to clusters smaller than a crossover lengthscale, beyond which the effects of disorder become relevant.
@@
  • Information about the central charge of the TLL can be extracted from the low-temperature behaviour of concurrence. We conjecture that this analysis can be extended to the effective central charge of the RS state, and show how this approach could be used to distinguish TLL from RS in the presence of quenched disorder.
  • The temperature–dependence of multi-partite entanglement, as witnessed by quantum Fisher information (QFI), provides a robust means of probing the effect of disorder on quantum spin systems. In the case of quantum spin chains, this provides a clear distinction between the TLL and RS states which can be accessed in experiment.
For our investigations, we employed exact diagonalization (ED) and the thermal pure quantum (TPQ) methods [69, 70, 71, 72]. Morevoer, throughout this Article we work in the units where $\hbar=1$ and $k_{B}=1$.
@@
In Sec. III, we detail our notation for evaluating averages. Sec. IV reviews key results related to the two states under investigation: the TLL and RS states. Sec. V introduces the model, disorder, and parameters used to investigate the entanglement structure of these states. Following which, in Sec. VI we present our zero temperature results for concurrence, RT and QFI. Sec. VII explores the finite temperature behavior of concurrence and equal time structure factor as witnesses. Here we address the subtleties involved with disorder averaging when dealing with converx roof measures like concurrence. We also discuss how to access the central charge for these states using concurrence and demonstrate the utility of using a multi-partite entanglement witness like equal-time structure factor as a probe for observing disorder effects.
In Sec. VIII, we discuss at length the nature of entanglement in the RS state, explaining the observed multi-partite entanglement and its origins. Finally, we conclude our findings in Sec. IX.
@@ II Accessing Entanglement in many body quantum systems
Entanglement in the modern perspective is thought of as a resource, one that can be used for quantum computation and information processing [73]. To this end, an important goal within quantum information is to quantify how entangled a state is? Unfortunately doing so is quite challenging and aside from a few special cases there is no solution as of yet. Even figuring out given a density matrix whether a state is separable or entangled, computationally, turns out to be very complex: in fact it is an NP-hard problem [74].
@@
In what we follows, we briefly review the concepts and definitions needed to understand this Article. We start by introducing the concept of entanglement, and then then introduce three entanglement witness-based measures: concurrence [66, 67], tangle [68], and quantum Fisher information (QFI) [40, 48]. The advantage of using these measures is that they can be measured from spin correlations and structure factors which are readily accessible by techniques like INS.
@@ II.1 Entanglement essentials
Having introduced the necessary terminology, we are in a good position to introduce the entanglement measures. We focus our investigation towards those suited for $S=1/2$ paricles. There do exists generalizations of the measures discussed for arbitrary spin quanta but we do not discuss them here. An exception is the QFI that we discuss in Sec. II.4
@@ II.2 Concurrence
Concurrence quantifies the pairwise entanglement between two $S=1/2$ particles. This could be in a system of only two or many $S=1/2$ particles. For a mixed state $\rho_{ij}$ shared between them where $i$ and $j$ are the indices pointing to the particles in question, it was shown [66, 67] to have the following relationship
@@
$\tilde{\rho}_{ij}=(\sigma_{y}\otimes\sigma_{y})\rho^{*}_{ij}(\sigma_{y}\otimes\sigma_{y})\;,$ (12)
@@ II.3 Residual Tangle
Residual tangle (RT) [68, 79] is a measure that accesses entanglement beyond the pairwise setting, in other words, a multi-partite measure. It is defined as the difference between one-tangle and the sum of pairwise entanglement measured via the two-tangle,
@@
is the one-tangle characterizing the entanglement of a $S=1/2$ particle indexed by $i$ with the rest of the system.
Strictly, $\tau_{1,i}$ serves as a measure of entanglement only for pure states. By extension, the same is true of $\tau_{3,i}$. And as a consequence, the interpretation of RT as an entanglement measure becomes inapplicable at finite temperatures. Moreover, although a nonzero RT suggests the presence of multi-partite entanglement, a zero RT does not necessarily indicate its absence.
@@ II.4 Quantum Fisher Information
from the dissipative part of dynamic susceptibility [48], $\chi^{\prime\prime}(\beta,\omega)$ defined as,
@@ III Notation
Figure 1: Comparison of the nearest and beyond nearest neighbor concurrence distribution in Tomonaga Luttinger liquid (TLL) and random singlet (RS) state. (a) Concurrence distribution between neighbors $n$ ranging from $1$ to $10$. (b) Concurrence distribution for $n>2$ neighbors. Here the bars indicate the height of the data values. As evident, the distribution of concurrence is localized only amongst the nearest neighbors for the TLL state while it is spread to further neighbors in the RS state. The plots here were obtained for a $N=30$ spin chain using Eq. (13) for the TLL, followed by an average of the sites. For the RS state, we evaluate Eq. (37) where $\Gamma$ is concurrence. For ease of viewing, the error bars have been suppressed.
@@ IV Overview of Heisenberg Spin chains
In this section we will briefly summarize necessary background pertaining to the Tomonaga Luttinger Liquid (TLL) and random singlet (RS) state in $d=1$. In this article, by TLL we refer to the long length scale theory of the Heisenberg (i.e. XXX) spin chain characterized by the Luttinger parameter, $K=1/2$ [82]. On the other hand, we interpret the strong disorder renormalization group (SDRG) [63] as the long length-scale theory of the RS state.
@@
In this case, the model realizes a $SU(2)$ symmetric ground state, whose long wavelength behavior provides an example of a TLL, with correlation function [83, 84]
@@
showcasong a power-law decay with exponent
@@
We now turn to the case where the couplings $J_{i}$ in Eq. (38) are positive, and randomly drawn from some probability distribution $P(J)$. In this case one realizes a RS state whose long-wavelength is well characterized by the SDRG framework [63]. Disorder-averaged correlation functions decay alegebraically
@@
independent of the choice of $P(J)$. Recently, however, logarithmic corrections to the above disorder averaged correlations have also been reported [85].
@@ V Setting up the problem
In this Section we make concrete our problem. Our goal here is to analyze what the previously introduced measures inform about the nature of entanglement of the TLL and RS states, respectively and whether one can use them in a meaningful way to distinguish between the two states. We investigate the antiferromagnetically interacting Heisenberg Hamiltonian in Eq. (38), with interactions drawn from a box distribution
@@
We make use of exact diagonalization for our analysis and use the quantum typicality method for finite temperature investigations [71, 72]. Throughout our analysis we restrict ourselves to spin chains of even length with periodic boundary conditions. For $T=0$ investigations we focussed on $M=1001$ disorder realizatons for $N$ ranging from $16$ to $30$. For $T\neq 0$, we consider $M=200$ disorder realizations for $N=16$ and $24$. For $N=20$, we consider $M=100$ disorder realizations. In addition, we have taken into account $100$ thermal pure quantum (TPQ) state samples for all three system sizes.
@@ VI Zero Temperature results
In what follows we will discuss the entanglement structure of TLL and RS states as characterized by concurrence, RT and QFI at zero temperature.
@@ VI.1 Nearest neighbor and beyond
We start by looking at the distribution of pairwise entanglement in the two states. In Fig. 1, we explore how the concurrence is distributed on the spin chain. For the RS state, we calculate the disorder-averaged concurrence i.e., we evaluate the $g^{zz}_{ij}$ correlation function for each realization and then from it obtain the corresponding concurrence using Eq. (13). Thereafter we average over the concurrence per disorder realization as per Eq. (37). For the TLL, we evaluate the sum in Eq. (37) without the disorder-average. As evident in Fig. 1, the TLL only has nearest neighbor pairwise entanglement [86, 87] while the RS also has a finite contribution coming from further neighbors as can be seen in Fig. 2 for chosen disorder realizations. However, relative to the contribution from the nearest neighbors it is quite small and in fact from the point of view of INS probes, is not observable. We explain this later in Sec. VII.1.
@@
Figure 2: Visually demonstrating the difference in the concurrence distribution between the random singlet (RS) and Tomonga Luttinger liquid (TLL) state respectively. (a) Concurrence distribution of the $147^{\text{th}}$ disorder realization. (b) Concurrence distribution of the $1000^{\text{th}}$ disorder realization. (c) Concurrence distribution of the TLL. While TLL only has nearest neighbor concurrence, the RS state has a non-zero concurrence for odd neighbors at all length scales. The plots here were obtained for a $N=30$ spin chain.

### v3

@@ Abstract
There are no clean samples in nature. Therefore, when we come to discuss the entanglement properties of quantum materials, the effects of disorder must be taken into account. This question is of particular interest for high-entangled phases, such as quantum spin liquids, which lie outside the Landau paradigm for classifying phases of matter. In this work, we explore what experimentally-accessible measures, in the form of concurrence, residual tangle and quantum Fisher information, can teach us about the entanglement in the presence of disorder. As a representative example, we consider the Tomonaga-Luttinger liquids (TLL) and disorder-driven random singlet (RS) phases found in antiferromagnetic quantum spin chains. Using quantum Fisher information and residual tangle, we demonstrate that both TLL and RS phases exhibit multi-partite entanglement. In the case of the RS phase, we attribute this to entanglement localized below a crossover length scale. We further show that the order of disorder average matters in calculating measures like concurrence, and that this can lead to false inferences when interpreting experiment. Nonetheless, correctly interpreted, these witnesses provide useful information about the effects of disorder. We explore how information about the central charge of the TLL can be extracted from the low-temperature behavior of concurrence, and conjecture that this analysis can be extended to the effective central charge of the RS phase. Finally, we establish how RS and TLL phases can be distinguished through the growth of multi-partite entanglement, as witnessed by the equal-time structure factor. These results establish that, used carefully, experiments based on entanglement witnesses can provide important information about quantum spin systems in the presence of disorder.
@@ I Introduction
How does disorder affect entanglement within quantum materials? Can we quantify such effects experimentally? In nature, all materials possess a certain degree of disorder. Thus, when trying to understand any phase of matter, the role of disorder is an important factor to take into consideration. Disorder or randomness is especially problematic when trying to identify quantum spin liquid (QSL) phases within candidate materials. This question arises in the context of many different materials, including $\kappa$-ET2Cu2(CN)3 [1, 2, 3, 4, 5, 6, 7, 8], Ba6Y2Rh2Ti2O17-δ [9, 10] , ZnCu3(OD)6Cl2 (Herbertsmithite) [11, 12, 13, 14, 15, 16, 7, 17, 18, 19, 20], YbMgGaO4 [21, 22, 23, 24], H3LiIr2O6 [25, 26, 27], etc. Here the challenge lies in being able to distinguish quantum fluctuations from disorder-induced ones. This puts into question whether it is the QSL or some disorder induced phase realized within these materials?
Entanglement is a valuable resource to identify phases such as QSLs [28]. Intuitively, one expects disorder, which localizes correlations, to have an effect on entanglement. The key issue, however, is whether this information can be accessed experimentally. Unlike cold atom systems where non local measures like (Rényi) entanglement entropy are accessible [29, 30], the same does not extend to condensed matter systems. Consequently, only limited information about the underlying phase can typically be accessed experimentally. It is thus important to determine if useful information about the entanglement structure can be obtained within experimental constraints. Can we do so by making use of experimentally-accessible entanglement measure [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61], for instance?
In this study, we explore how measures based on entanglement witnesses can be used to distinguish the different phases which arise in quantum spin systems with and without disorder. As a representative example, we consider the $S=1/2$ Heisenberg (XXX) spin chain with and without bond disorder. The competing phases in this case are the Tomonaga Luttinger liquid (TLL) [62] found in the absence of disorder, and the random singlet (RS) phase realised for arbitrarily weak bond disorder [63, 64, 65, 66]. We consider three different measures, namely concurrence [67, 68], tangle [69] and quantum Fisher information (QFI) [42, 50], all of which have previously been discussed in the context of experiment [55, 57, 59, 61], and use exact diagonalization to explore how each of these behaves in the TLL and RS phases.
From our analysis, we conclude:
  • Both the TLL and RS phases are multi-partite entangled. In the TLL, multi-partite entanglement is present at all length scales. However, in the RS phase, multi-partite entanglement is restricted to clusters smaller than a crossover length-scale, beyond which the effects of disorder become relevant.
@@
  • Information about the central charge of the TLL can be extracted from the low-temperature behaviour of concurrence. We conjecture that this analysis can be extended to the effective central charge of the RS phase, and show how this approach could be used to distinguish TLL from RS in the presence of quenched disorder.
  • The temperature–dependence of multi-partite entanglement, as witnessed by quantum Fisher information (QFI), provides a robust means of probing the effect of disorder on quantum spin systems. In the case of quantum spin chains, this provides a clear distinction between the TLL and RS phases which can be accessed in experiment.
For our investigations, we employed exact diagonalization (ED) and the thermal pure quantum (TPQ) methods [70, 71, 72, 73]. Moreover, throughout this Article we work in the units where $\hbar=1$ and $k_{B}=1$.
For completeness, we note that concurrence at zero temperature has previously been discussed in the context of disordered spin chains [74, 75]. An interesting alternative approach distinguishing RS from the TLL phase employing Bell’s inequality was discussed in [74].
@@
In Sec. III, we detail our notation for evaluating averages. Sec. IV reviews key results related to the two states under investigation: the TLL and RS phases. Sec. V introduces the model, disorder, and parameters used to investigate the entanglement structure of these phases. Following which, in Sec. VI we present our zero temperature results for concurrence, RT and QFI. Sec. VII explores the finite temperature behavior of concurrence and equal time structure factor as witnesses. Here we address the subtleties involved with disorder averaging when dealing with convex roof measures like concurrence. We also discuss how to access the central charge for these phases using concurrence and demonstrate the utility of using a multi-partite entanglement witness like equal-time structure factor as a probe for observing disorder effects.
In Sec. VIII, we discuss at length the nature of entanglement in the RS phase, explaining the observed multi-partite entanglement and its origins. Finally, we conclude our findings in Sec. IX.
@@ II Accessing Entanglement in many body quantum systems
Entanglement in the modern perspective is thought of as a resource, one that can be used for quantum computation and information processing [76]. To this end, an important goal within quantum information is to quantify how entangled a state is? Unfortunately doing so is quite challenging and aside from a few special cases there is no solution as of yet. Even figuring out, given a density matrix, whether a state is separable or entangled, computationally, turns out to be very complex: in fact it is an NP-hard problem [77].
@@
In what we follows, we briefly review the concepts and definitions needed to understand this Article. We start by introducing the concept of entanglement, and then introduce three entanglement witness-based measures: concurrence [67, 68], tangle [69], and quantum Fisher information (QFI) [42, 50]. The advantage of using these measures is that they can be measured from spin correlations and structure factors which are readily accessible by techniques like inelastic Neutron scattering (INS).
@@ II.1 Entanglement essentials
Having introduced the necessary terminology, we are in a good position to introduce the entanglement measures. We focus our investigation towards those suited for $S=1/2$ particles. There do exist generalizations of the measures discussed for arbitrary spin quanta but we do not discuss them here. An exception is the QFI that we discuss in Sec. II.4
@@ II.2 Concurrence
Concurrence quantifies the pairwise entanglement between two $S=1/2$ particles, by quantifying the “entanglement of formation” [79], required to construct a given state. This could be in a system of only two or many $S=1/2$ particles. For a mixed state $\rho_{ij}$ shared between them where $i$ and $j$ are the indices pointing to the particles in question, it was shown [67, 68] to have the following relationship
@@
$\tilde{\rho}_{ij}=(\sigma^{y}\otimes\sigma^{y})\rho^{*}_{ij}(\sigma^{y}\otimes\sigma^{y})\;,$ (12)
@@
As a parting comment, we note that in the case where a pair of spins exist in a pure state, the definition of concurrence reduces to a simpler form [67]. However when considering states with multipartite entanglement, it is necessary to use the more general definition, Eq. (10), since the reduced density matrix for a pair of spins will not, in general, be a pure state. This point is addressed in greater detail Appendix A.
@@ II.3 Residual Tangle
Residual tangle (RT) [69, 83, 61] is a measure that accesses entanglement beyond the pairwise setting, in other words, a multi-partite measure. It is defined as the difference between one-tangle and the sum of pairwise entanglement measured via the two-tangle,
@@
is the one-tangle [69, 83, 61].
Strictly, $\tau_{3,i}$ is only well defined for pure states [61]. Consequently, the interpretation of RT as an entanglement measure becomes inapplicable at finite temperatures. Moreover, although a nonzero RT suggests the presence of multi-partite entanglement, a zero RT does not necessarily indicate its absence.
@@ II.4 Quantum Fisher Information
from $\chi^{\prime\prime}(\beta,\omega)$, the imaginary, dissipative part of dynamic susceptibility [50], $\chi(\beta,\omega)$ defined as,
@@ III Notation
Figure 1: Comparison of distribution of concurrence as a function of separation of spins within the Tomonaga Luttinger liquid (TLL) and random singlet (RS) phases. (a) Distribution of concurrence $C_{ij}$ on $n^{th}$ neighbour bonds, for $n=1\ldots 10$. Within the TLL (orange bars), concurrence vanishes identically for all bonds with $n>1$. In contrast, within the RS phase (green bars), concurrence takes on a finite value for all odd values of $n$. (b) Detail of $C_{ij}$ for $n>2$, showing slow decay of concurrence within the RS phase. All results are shown are taken from exact diagonalization of an $N=30$ spin chain, with $C_{ij}$ defined through Eq. (13), and disorder averages calculated for $M=1001$ instances, using Eq. (37), with $\Gamma=C_{ij}$. For ease of viewing, error bars have been suppressed.
@@ IV Overview of Heisenberg Spin chains
In this section we will briefly summarize necessary background pertaining to the Tomonaga Luttinger Liquid (TLL) and random singlet (RS) phase in $d=1$. In this article, by TLL we refer to the long length scale theory of the Heisenberg (i.e. XXX) spin chain characterized by the Luttinger parameter, $K=1/2$ [86]. On the other hand, we interpret the strong disorder renormalization group (SDRG) [65] as the long length-scale theory of the RS phase.
@@
In this case, the model realizes a $SU(2)$ symmetric ground state, whose long wavelength behavior provides an example of a TLL phase, with correlation function [87, 88]
@@
showcasing a power-law decay with exponent
@@
We now turn to the case where the couplings $J_{i}$ in Eq. (38) are positive, and randomly drawn from some probability distribution $P(J)$. In this case one realizes a RS phase whose long-wavelength is well characterized by the SDRG framework [65]. Disorder-averaged correlation functions decay algebraically
@@
independent of the choice of $P(J)$. We note that logarithmic corrections to power-law decay, Eq. (42), have been reported [89]. However, their validity has been questioned in subsequent studies [90, 91].
@@ V Setting up the problem
In this Section we make concrete our problem. Our goal here is to analyze what the previously introduced measures inform about the nature of entanglement of the TLL and RS phases, respectively and whether one can use them in a meaningful way to distinguish between the two phases. We investigate the antiferromagnetically interacting Heisenberg Hamiltonian in Eq. (38), with interactions drawn from a box distribution
@@
We make use of exact diagonalization for our analysis and use the quantum typicality method for finite temperature investigations [72, 73]. Throughout our analysis we restrict ourselves to spin chains of even length with periodic boundary conditions. For $T=0$ investigations we focused on $M=1001$ disorder realizatons for $N$ ranging from $16$ to $30$. For $T\neq 0$, we consider $M=200$ disorder realizations for $N=16$ and $24$. For $N=20$, we consider $M=100$ disorder realizations. In addition, we have taken into account $100$ thermal pure quantum (TPQ) state samples for all three system sizes.
@@ VI Zero Temperature results
In what follows we will discuss the entanglement structure of TLL and RS phases as characterized by concurrence, RT and QFI at zero temperature.
@@ VI.1 Nearest neighbor and beyond
We start by looking at the distribution of pairwise entanglement in the two phases. In Fig. 1, we explore how the concurrence is distributed on the spin chain. For the RS phase, we calculate the disorder-averaged concurrence i.e., we evaluate the $g^{zz}_{ij}$ correlation function for each realization and then from it obtain the corresponding concurrence using Eq. (13). Thereafter we average over the concurrence per disorder realization as per Eq. (37). For the TLL, we evaluate the sum in Eq. (37) without the disorder-average.
As evident in Fig. 1, the TLL only has nearest neighbor pairwise entanglement [92, 93] while the RS also has a finite contribution coming from further neighbors as can be seen in Fig. 2 for chosen disorder realizations. The concurrence of the RS phase vanishes at even distance, $n=\{2,4,6,\ldots\}$, while showing a slow decay at odd distance $n=\{1,3,5,\ldots\}$, consistent with earlier calculations for disordered XY spin chains [74]. This echos the patterns found in spin correlations at long distances where the decay is controlled by rare, strong singlet bonds [65, 94]. None the less, it would be a mistake to identify the concurrence of the RS singlet phase identically with its correlations, a point addressed in Appendix A.
We note that, relative to the contribution from the nearest neighbors the concurrence observed at longer length scales is quite small, and in fact from the point of view of INS probes, is not observable. We explain this later in Sec. VII.1.
@@
Figure 2: Visually demonstrating the difference in the concurrence distribution between the random singlet (RS) and Tomonaga Luttinger liquid (TLL) phase respectively. (a) Concurrence distribution of the $147^{\text{th}}$ disorder realization. (b) Concurrence distribution of the $1000^{\text{th}}$ disorder realization. (c) Concurrence distribution of the TLL. While TLL only has nearest neighbor concurrence, the RS phase has a non-zero concurrence for odd neighbors at all length scales. The plots here were obtained for a $N=30$ spin chain.